OpenStudy (anonymous):

.Ashopkeeperwantstoearn12% profitonanitemafter giving20% discount tothe customer. Bywhatpercentage shouldhe increasehismarkedprice toarriveat the label price?

OpenStudy (anonymous):

profit=12% selling price =80x/100 marked price =x

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

so he gets 80% of marketprice

ganeshie8 (ganeshie8):

(80% of market price ) - (cost price) = (12% of cost price)

ganeshie8 (ganeshie8):

0.8*m - c = 0.12*c 0.8*m = 1.12c m = (1.12/0.8)c = 1.4c

ganeshie8 (ganeshie8):

that means he needs to increase the cost price by 40% to achieve the desired profit

OpenStudy (anonymous):

yes absolutely

OpenStudy (anonymous):

i need to solve some logical question do u know that

OpenStudy (anonymous):

@perl